# Multi-scale direction-aware SAR object detection network via global information fusion

### v4

@@ Abstract
Deep learning has driven significant progress in object detection using Synthetic Aperture Radar (SAR) imagery. Existing methods, while achieving promising results, often struggle to effectively integrate local and global information, particularly direction-aware features. This paper proposes SAR-Net, a novel framework specifically designed for global fusion of direction-aware information in SAR object detection. SAR-Net leverages two key innovations: the Unity Compensation Mechanism (UCM) and the Direction-aware Attention Module (DAM). UCM facilitates the establishment of complementary relationships among features across different scales, enabling efficient global information fusion and transmission. Additionally, DAM, through bidirectional attention polymerization, captures direction-aware information, effectively eliminating background interference. Extensive experiments demonstrate the effectiveness of SAR-Net, achieving state-of-the-art results on aircraft (SAR-AIRcraft-1.0) and ship datasets (SSDD, HRSID), confirming its generalization capability and robustness.
keywords:
Deep learning , SAR object detection , Global fusion , Direction-aware attention
[…]
{graphicalabstract}
{highlights}
A SAR object detector to tackle information fusion and direction-aware problems.
Unity Compensation Mechanism (UCM) and Direction-aware Attention Module (DAM).
UCM unites multi-scale features and compensates for the original features.
DAM obtains positional and directional information in complex background.
We achieve SOTA accuracy among SAR-AIRcraft-1.0, SSDD and HRSID datasets.
1 Introduction
Object detection plays a pivotal role in remote sensing, with extensive applications in military operations, security surveillance, disaster assessment, and environmental monitoring [1]. Synthetic Aperture Radar (SAR) object detection, a key technique in remote sensing, offers notable benefits in super-resolution imaging, hidden object detection, and all-weather surveillance, particularly in identifying aircraft and ships.
In recent years, with the rapid development of deep learning, an increasing number of researchers have turned their attention to design object detection methods using Convolutional Neural Networks (CNNs), yielding notable results. However, the unique challenges posed by SAR images, characterized by their multi-scale and directional nature, coupled with abundant background details, complicate object detection tasks. These factors particularly hinder the detection of small objects amidst background noise, making the design of effective multi-scale object detection networks for SAR images a complex task. Consequently, developing strategies to accurately determine the direction and position of aircraft or ships in SAR images, despite background interference, stands as a critical challenge in this field.
@@ 1 Introduction
2 Related work
2.1 SAR object detection
Object detection in SAR images has gained significant attention in recent years. Traditional detection algorithms, which follow the Constant False Alarm Rate (CFAR) principles [9, 10, 11], operated based on a statistical model. In prior works [12], researchers aimed to segregate objects from the background to emphasize disparities and streamline detection processes. Some other works [13] were devoted to solving the problem of multi-scale object detection in high-resolution images. Despite these achievements, traditional methods still face constraints in detecting small objects and heavily depend on manual design, posing challenges for robustness and limiting their applicability to different SAR datasets. Hence, there is a critical demand for more adaptable and automated approaches in SAR object detection.
Owing to the inherent capacity for autonomous feature learning in deep learning, researchers have seamlessly integrated object detection in SAR images with deep learning methodologies [14]. Shifting focus to SAR ship detection and identification, Zhang et al. [15] proposes a joint unsupervised super-resolution and SAR object detection network. This network augments object details while suppressing noise, providing valuable guidance to the detection network. In the realm of ship detection and recognition, Guo et al. [14] developed three modules for feature refinement and fusion, while utilizing a head module to reduce the impact of background noise. Furthermore, Li et al. [2] utilized the FPN to create spatial multi-scale features, specifically for ship objects, through a top-down hierarchical approach. In a related endeavor, Li et al. [5] introduced the AFF module to strategically integrate local shallow features into adjacent layers. This self-decision approach enables the flexible integration of different features, enhancing the model’s ability to extract object features while also providing a certain level of robustness. Addressing the specific challenge of large-scale SAR images, Yang et al. [16] holistically considers both pre-processing and post-processing in algorithm optimization, leading to an overall improvement in the effectiveness of object detection in large-scale SAR images. The comprehensive strategy employed by these studies underscores the effectiveness of deep learning paradigms in advancing the state-of-the-art in SAR object detection and recognition.
@@ 2.1 SAR object detection
2.2 Multi-scale object detection
In the realm of deep learning, the adoption of image pyramids initially emerged as a solution for tackling multi-scale challenges in object detection, driven by its intuitive appeal. Notable approaches distinguished themselves by selecting specific scales for each resolution during multi-scale training. However, it’s essential to note that these methods significantly increase inference time, creating a trade-off between accuracy and efficiency. Another approach is to build a feature pyramid inside the network, providing an approximate image pyramid and reducing computational costs. This enables the detection of multi-scale objects at each feature layer but can lead to decreased accuracy in detecting small objects. The issue arises from the low-level features themselves lacking semantic information, which is particularly crucial for small object detection.
To address the semantic limitations in low-level features, the Feature Pyramid Network (FPN) [3] is introduced. FPN offers an efficient structural design for merging multi-scale features via skip connections and local information transfer. Researchers at FSAF [17] identify a flaw in FPN’s heuristic-guided feature selection, highlighting the need for improvement. In contrast, FSAF focuses attention on regions of interest within each hierarchical feature level, aiming to extract features from a single level, but ignores valuable information between different hierarchical levels. The SA-FPN [18], amalgamating Top-Down style FPN and Bottom-Up style FPN, combined the strengths of both, emerging as a more accurate module for perceiving scale variations. While previous methods integrate information at all levels through common operations like element-wise max or sum, these fusion mechanisms do not consider semantic differences among feature layers. EfficientDet [19] addressed this limitation by introducing a repeatable module (BiFPN) to enhance information fusion efficiency between different levels. Recognizing that the feature maps in a pyramid predominantly consist of single-level features, M2Det [20] provided insights into this structure. Moreover, [21] extended FPN with the Asymptotic Feature Pyramid Network (AFPN) to facilitate interaction across non-adjacent layers. To overcome FPN’s limitations in detecting large objects, [22] proposed a parallel FPN structure for object detection with bi-directional fusion.
@@ 3.1 Overall framework of SAR-Net
3.2 Unity compensation mechanism
@@
3.2.1 Multi-scale alignment module
@@
3.2.2 Multi-level fusion module
@@
3.2.3 Multi-feature embedding module
Inspired by the fusion architecture with two branches, to embed global information more effectively at different levels, we employ attention operations for information fusion, as illustrated in Fig. 5. In particular, we feed local information (pertaining to features at the operating level) and globally embedded information (generated by MFM) into two separate 1D convolutions for feature transformation. The Sigmoid function is utilized to generate attention weights from global information, which are then applied to weight the local information. Additionally, for further compensation, a residual connection is employed. This is aimed at achieving deep fusion and enhancement of global and local information. Finally, the DA Block is used for further feature extraction and fusion.
3.3 Direction-aware attention module
@@
3.3.1 Direction-aware generation
Due to the dominance of background in SAR image content and the presence of interference areas such as docks, airports, and terminals, directly locating and detecting objects becomes challenging. Additionally, SAR images often exhibit the edge contours of objects, reflecting distinct directional features, while extracting detailed information within the objects is challenging. Therefore, leveraging the stability and distinctiveness of object direction information can effectively reduce the difficulty of localization.
@@ 3.3.1 Direction-aware generation
3.3.2 Channel attention embedding
@@
3.4 Loss function
The overall loss of our network consists of two components: the loss of object classification $L_{cls}$ and the loss related to bounding box regression $L_{reg}$. It can be expressed as:
@@ 3.4 Loss function
we adopt the VariFocal Loss (VFL) as the classification loss $L_{cls}$. VFL balances the learning signals between positive and negative samples by considering their varying levels of importance. The bounding box regression loss $L_{reg}$ comprises IOU loss and probability loss, as shown below:
@@
here, $\lambda_{iou}$ and $\lambda_{dfl}$ are hyperparameters. We choose the advanced SIOU as the IOU loss $L_{iou}$. Due to the consideration of ambiguity and uncertainty in the data without introducing additional strong priors, the Distribution Focal Loss (DFL) contributes to improving box localization accuracy. Additionally, we add a necessary loss $L_{dfl}$, which accounts for the bounding box regression loss.
4 Experiments and analysis
@@
4.1 Dataset and implementation details
Table 1: The total of instance objects of different categories on the SAR-AIRcraft-1.0.
@@ 4.1 Dataset and implementation details
4.1.1 SAR-AIRcraft-1.0
Images acquired from the Gaofen-3 satellite are screened to form the SAR-AIRcraft-1.0 dataset. The polarization method is single polarization, with a spatial resolution of $1m$, and utilizing the spotlight imaging mode. Considering the size of the airport and the number of parked aircraft, the dataset mainly uses image data of three civil airports: Shanghai Hongqiao Airport, Beijing Capital Airport, and Taiwan Taoyuan Airport, including a total of $800\times 800$, $1000\times 1000$, $1200\times 1200$ and $1500\times 1500$. 4 different sizes, a total of 4,368 pictures and 16,463 aircraft object instances. The specific categories of aircraft include A220, A320/321, A330, ARJ21, Boeing737, Boeing787, other. The total of instance objects of different categories is shown in Table 1, where other represents aircraft instances that do not belong to the other 6 categories. The size distribution of the aircraft object slices in the dataset spans a wide range. As shown in Fig. 7 (a), some objects are smaller than $50\times 50$, and some aircraft objects are larger than $100\times 100$. The overall object size is multi-scale. In addition, in this paper, we use the default dataset partitioning by [25] and resize all images to $800\times 800$ pixels. To ensure fairness, we adopt the above settings when compared with other methods.
4.1.2 SSDD
4.1.3 HRSID
4.2 Accuracy metrics
$Precision=\frac{TP}{TP+FP},$ (14)
$Recall=\frac{TP}{TP+FN},$ (15)
[…]
4.3 Ablation study
Table 2: Ablation study of three modules on unity compensation mechanism with MAM using concatenation operation.
@@ 4.3 Ablation study
Table 3: Ablation study of three modules on unity compensation mechanism with MAM using addition operation.
@@
4.3.1 Impact of unity compensation mechanism
To study the effectiveness of each module in the unity compensation mechanism and verify the correctness of uniting deep and shallow layers’ global information, we executed a lot of ablation experiments on four of the modules and measured the performance with mAP50 and F1-score, a common indicator in object detection. In addition, Parameter sizes and GFLOPs are also shown in Table 2 and Table 3. We use YOLOv6n [23] as the baseline for ablation experiments to embed the modules of our unity compensation mechanism one by one. The feature alignment process is included in all experiments except the baseline. In Table 2, we conducted MAM using concatenation operation, where S-MFM represents Shallow-MFM and D represents Deep-MFM. It can be seen that only when the MAM module is used, the accuracy is lower than the baseline. This is because it integrates semantic information and detailed texture information from different layers, but it lacks further fusion of object-specific information. In addition, other experiments have achieved performance improvements without a substantial increase in computational overhead, which proves the efficacy of our global information fusion idea. In particular, when the MEM module is not used, we use simple addition to compensate for the original features.
Figure 8: Layercam visualization of different modules on Unity Compensation Mechanism.
To explore a new alignment pattern, we conducted MAM using addition operation. It is worth noting that both concatenation and addition operations require aligning features from each layer to have the same dimensions. As can be seen from Table 3, the overall trends of the other modules are similar to those in Table 2. However, due to the addition operation combining rich texture and semantic information directly from different layers, it made the information redundant and resulted in difficulties in subsequent fusion compensation operations, leading to information loss and performance degradation. Additionally, features from different layers have varying channel numbers, requiring channel alignment during the addition operation, which increases parameter and computation complexity.
Table 4: Research on the effectiveness and optimal embedding locations of direction-aware attention module.
@@ 4.3.1 Impact of unity compensation mechanism
4.3.2 Impact of direction-aware attention module
To prove the effectiveness of the proposed direction-aware attention module, we executed a series of detailed experiments, as shown in Table 4. We split the attention modules in two directions and embedded them in the YOLOv6n backbone separately, both received performance improvements. To explore the roles of the two modules within DAM, we conducted experiments by using only one module at a time on the baseline. Here, w/o CAE represents without Channel Attention Embedding, and w/o DG represents without Direction-aware Generation. Both experiments showed improvement compared to the baseline. However, due to the absence of crucial bidirectional feature vector extraction, DAM w/o DG showed minimal performance enhancement, merely reallocating attention among channels. Equally noteworthy, after generating the direction-aware attention map, omitting the embedding of the attention map among channels also resulted in performance degradation. Therefore, the collaboration of these two modules enables the model to concentrate more effectively on the directional areas of objects.
Table 5: Collaborative research within the framework, UC represents unity compensation.
@@ 4.3.2 Impact of direction-aware attention module
4.3.3 Collaboration of the framework
The effectiveness of the components that make up SAR-Net has been demonstrated, and to further explore the synergy between them, we combine them and find the desired results. For the unity compensation mechanism, the backbone that can extract direction-aware information gives full play to the advantages of global fusion and enriches the information of multi-scale features. For the DAM-embedded backbone, the unity compensation mechanism retains detailed information and enhances the efficiency of information fusion and transmission. As shown in Table 5, after combining the two, their collaborative complementarity is fully utilized and performance is further improved by 0.4%.
Table 6: Comparison with different FPN variants.
@@ 4.3.3 Collaboration of the framework
4.3.4 Comparison of different FPN architecture
We choose the classic FPN structure and its latest variants AFPN [21] and RepBi-PAN [23] for comparison. We conduct the experiments based on YOLOv6n and the results are illustrated in Table 6. A significant accuracy disparity exists despite the parameters being similar. As we discuss in the section 1, sequential local fusion will lead to information loss, especially for SAR images with large object scale variants. Top-down and bottom-up structures will lose a lot of detailed information and semantic information before detection. Global information fusion and compensation preserve the object’s multi-scale information to the maximum extent, thereby enabling the unity compensation mechanism to deliver optimal performance.
Table 7: Comparison with other methods on SAR-AIRcraft-1.0. The accuracy of each category is represented by mAP with an intersection over the union threshold of 0.5.

### v5

# SAR-Net: Multi-scale Direction-aware SAR Network via Global Information Fusion

@@ Abstract
Deep learning has driven significant progress in object detection using Synthetic Aperture Radar (SAR) imagery. Existing methods, while achieving promising results, often struggle to effectively integrate local and global information, particularly direction-aware features. This paper proposes SAR-Net, a novel framework specifically designed for global fusion of direction-aware information in SAR object detection. SAR-Net leverages two key innovations: the Unity Compensation Mechanism (UCM) and the Direction-aware Attention Module (DAM). UCM facilitates the establishment of complementary relationships among features across different scales, enabling efficient global information fusion. Among them, Multi-scale Alignment Module (MAM) and distinct Multi-level Fusion Module (MFM) enhance feature integration by capturing both texture detail and semantic information. Then, Multi-feature Embedding Module (MEM) feeds back global features into the primary branches, further improving information transmission. Additionally, DAM, through bidirectional attention polymerization, captures direction-aware information, effectively eliminating background interference. Extensive experiments demonstrate the effectiveness of SAR-Net, achieving state-of-the-art results on aircraft (SAR-AIRcraft-1.0) and ship datasets (SSDD, HRSID), confirming its generalization capability and robustness.
Index Terms:
Deep learning, Remote Sensing, SAR object detection, global fusion, direction-aware attention.
I Introduction
OBJECT detection plays a pivotal role in remote sensing, with extensive applications in military operations, security surveillance, disaster assessment, and environmental monitoring[1, 2]. Synthetic Aperture Radar (SAR) object detection, a key technique in remote sensing, offers notable benefits in super-resolution imaging, hidden object detection, and all-weather surveillance, particularly in identifying aircraft and ships.
In recent years, with the rapid development of deep learning, an increasing number of researchers have turned their attention to design object detection methods using Convolutional Neural Networks (CNNs), yielding notable results [3, 4, 5, 6, 7, 8, 9]. However, the unique challenges posed by SAR images, characterized by their multi-scale and directional nature, coupled with abundant background details, complicate object detection tasks. These factors particularly hinder the detection of small objects amidst background noise, making the design of effective multi-scale object detection networks for SAR images a complex task. Consequently, developing strategies to accurately determine the direction and position of aircraft or ships in SAR images, despite background interference, stands as a critical challenge in this field.
@@ I Introduction
II Related Work
II-A SAR Object Detection
Object detection in SAR images has gained significant attention in recent years. Traditional detection algorithms, which follow the Constant False Alarm Rate (CFAR) principles [17], operated based on a statistical model. In prior works [18], [19], researchers aimed to segregate objects from the background to emphasize disparities and streamline detection processes. Some other works [20, 21, 22, 23] were devoted to solving the problem of multi-scale object detection in high-resolution images. Despite these achievements, traditional methods still face constraints in detecting small objects and heavily depend on manual design, posing challenges for robustness and limiting their applicability to different SAR datasets. Hence, there is a critical demand for more adaptable and automated approaches in SAR object detection.
Owing to the inherent capacity for autonomous feature learning in deep learning, researchers have seamlessly integrated object detection in SAR images with deep learning methodologies[24, 25, 26, 27, 28, 29]. Shifting focus to SAR ship detection and identification, Zhang et al. [30] proposes a joint unsupervised super-resolution and SAR object detection network. This network augments object details while suppressing noise, providing valuable guidance to the detection network. In the realm of SAR aircraft detection and recognition, Zhao et al. [31] introduces an attention module, strategically fusing refined low-level texture features with high-level semantic features, which enhances aircraft detection rates by leveraging complementary information. Furthermore, Zhao et al. [32] devises a multi-branch dilated convolution feature pyramid method. This innovative approach minimizes redundancy while accentuating critical aircraft features through the establishment of dense connections. In a related endeavor, Zhao et al. [33] put forth a multitask learning framework for object detection (MLDet) tailored for ship detection in SAR imagery. This end-to-end framework encompasses several tasks and had a good effect. The collaborative execution of these tasks contributes to an enhanced detection performance. Addressing the specific challenge of large-scale SAR images, Yang et al. [34] holistically considers both pre-processing and post-processing in algorithm optimization, leading to an overall improvement in the effectiveness of object detection in large-scale SAR images. The comprehensive strategy employed by these studies underscores the effectiveness of deep learning paradigms in advancing the state-of-the-art in SAR object detection and recognition.
@@ II-A SAR Object Detection
II-B Multi-scale Object Detection
In the realm of deep learning, the adoption of image pyramids initially emerged as a solution for tackling multi-scale challenges in object detection, driven by its intuitive appeal [35, 36, 37, 38, 39]. Notable approaches like SNIP [40] and SNIPER [41] distinguished themselves by selecting specific scales for each resolution during multi-scale training. However, it’s essential to note that these methods significantly increase inference time, creating a trade-off between accuracy and efficiency. Another approach is to build a feature pyramid inside the network, providing an approximate image pyramid and reducing computational costs, as seen in SSD [42]. This enables the detection of multi-scale objects at each feature layer but can lead to decreased accuracy in detecting small objects. The issue arises from the low-level features themselves lacking semantic information, which is particularly crucial for small object detection.
To address the semantic limitations in low-level features, the Feature Pyramid Network (FPN) [11] is introduced. FPN offers an efficient structural design for merging multi-scale features via skip connections and local information transfer. Building upon FPN, the Path Aggregation Network (PANet) [43] introduces a bottom-up path to enhance information fusion between different levels. Researchers at FSAF [44] identify a flaw in FPN’s heuristic-guided feature selection, highlighting the need for improvement. In contrast, FSAF focuses attention on regions of interest within each hierarchical feature level, aiming to extract features from a single level, but ignores valuable information between different hierarchical levels. The SA-FPN [45], amalgamating Top-Down style FPN and Bottom-Up style FPN, combined the strengths of both, emerging as a more accurate module for perceiving scale variations. While previous methods integrate information at all levels through common operations like element-wise max or sum, these fusion mechanisms do not consider semantic differences among feature layers. EfficientDet [46] addressed this limitation by introducing a repeatable module (BiFPN) to enhance information fusion efficiency between different levels. YOLO-ASFF [47] employed an adaptive spatial feature fusion technique, learning weights during training to filter out inconsistencies. Recognizing that the feature maps in a pyramid predominantly consist of single-level features, M2Det [48] provided insights into this structure. Moreover, [49] extended FPN with the Asymptotic Feature Pyramid Network (AFPN) to facilitate interaction across non-adjacent layers. To overcome FPN’s limitations in detecting large objects, [50] proposed a parallel FPN structure for object detection with bi-directional fusion, while [51] introduced a refined FPN structure. SAFNet [52] introduced Adaptive Feature Fusion and Self-Enhanced Modules.
@@ III-A Overall Framework of SAR-Net
III-B Unity Compensation Mechanism
@@
III-B1 Multi-scale Alignment Module
@@
III-B2 Multi-level Fusion Module
@@
III-B3 Multi-feature Embedding Module
Inspired by the fusion architecture with two branches in [55], to embed global information more effectively at different levels, we employ attention operations for information fusion, as illustrated in Fig. 5. In particular, we feed local information (pertaining to features at the operating level) and globally embedded information (generated by MFM) into two separate 1D convolutions for feature transformation. The Sigmoid function is utilized to generate attention weights from global information, which are then applied to weight the local information. Additionally, for further compensation, a residual connection is employed. This is aimed at achieving deep fusion and enhancement of global and local information. Finally, the DA Block is used for further feature extraction and fusion.
III-C Direction-aware Attention Module
@@
III-C1 Direction-aware Generation
Due to the dominance of background in SAR image content and the presence of interference areas such as docks, airports, and terminals, directly locating and detecting objects becomes challenging. Additionally, SAR images often exhibit the edge contours of objects, reflecting distinct directional features, while extracting detailed information within the objects is challenging. Therefore, leveraging the stability and distinctiveness of object direction information can effectively reduce the difficulty of localization [56].
@@ III-C1 Direction-aware Generation
III-C2 Channel Attention Embedding
@@
III-D Loss Function
The overall loss of our network consists of two components [53]: the loss of object classification $L_{cls}$ and the loss related to bounding box regression $L_{reg}$. It can be expressed as:
@@ III-D Loss Function
we adopt the VariFocal Loss (VFL) [57] as the classification loss $L_{cls}$. VFL balances the learning signals between positive and negative samples by considering their varying levels of importance. The bounding box regression loss $L_{reg}$ comprises IOU loss and probability loss, as shown below:
@@
here, $\lambda_{iou}$ and $\lambda_{dfl}$ are hyperparameters. We choose the advanced SIOU [58] as the IOU loss $L_{iou}$. Due to the consideration of ambiguity and uncertainty in the data without introducing additional strong priors, the Distribution Focal Loss (DFL) [59] contributes to improving box localization accuracy. Additionally, we add a necessary loss $L_{dfl}$, which accounts for the bounding box regression loss.
IV Experiments and Analysis
@@
TABLE I: The Total of Instance Objects of Different Categories on the SAR-AIRcraft-1.0
@@
IV-A Dataset and Implementation Details
TABLE II: Ablation Study of Three Modules on Unity Compensation Mechanism with MAM Using Concatenation Operation
@@ IV-A Dataset and Implementation Details
IV-A1 SAR-AIRcraft-1.0
Images acquired from the Gaofen-3 satellite are screened to form the SAR-AIRcraft-1.0 dataset. The polarization method is single polarization, with a spatial resolution of $1m$, and utilizing the spotlight imaging mode. Considering the size of the airport and the number of parked aircraft, the dataset mainly uses image data of three civil airports: Shanghai Hongqiao Airport, Beijing Capital Airport, and Taiwan Taoyuan Airport, including a total of $800\times 800$, $1000\times 1000$, $1200\times 1200$ and $1500\times 1500$. 4 different sizes, a total of 4,368 pictures and 16,463 aircraft object instances. The specific categories of aircraft include A220, A320/321, A330, ARJ21, Boeing737, Boeing787, other. The total of instance objects of different categories is shown in Table I, where other represents aircraft instances that do not belong to the other 6 categories. The size distribution of the aircraft object slices in the dataset spans a wide range. As shown in Fig. 7 (a), some objects are smaller than $50\times 50$, and some aircraft objects are larger than $100\times 100$. The overall object size is multi-scale. In addition, in this paper, we use the default dataset partitioning by [60] and resize all images to $800\times 800$ pixels. To ensure fairness, we adopt the above settings when compared with other methods.
TABLE III: Ablation Study of Three Modules on Unity Compensation Mechanism with MAM Using Addition Operation
@@
Figure 8: Layercam [63] visualization of different modules on Unity Compensation Mechanism.
IV-A2 SSDD
TABLE IV: Research on the Effectiveness and Optimal Embedding Locations of Direction-aware Attention Module
@@ IV-A2 SSDD
IV-A3 HRSID
TABLE V: Collaborative Research Within the Framework, UC Represents Unity Compensation
@@ IV-A3 HRSID
TABLE VI: Comparison with Different FPN Variants
@@
IV-B Accuracy Metrics
$Precision=\frac{TP}{TP+FP},$ (14)
$Recall=\frac{TP}{TP+FN},$ (15)
[…]
TABLE VII: Comparison with Other Methods on SAR-AIRcraft-1.0. The Accuracy of Each category is Represented by mAP with an Intersection Over the Union Threshold of 0.5